# Fundamental role of Dirac supersingletons in particle theory

### v1

@@ 5 Conclusion
As shown in the famous Dyson’s paper [3], even from purely mathematical considerations (without any physics) it is clear that Poincare quantum symmetry is a special degenerate case of de Sitter quantum symmetries. As shown in [2, 4], in standard quantum theory (over the field of complex numbers), dS symmetry is more general than AdS one but in the framework of this symmetry it is not possible to describe neutral elementary particles, i.e., particles which are equivalent to the their antiparticles. In particular, even the photon cannot be a neutral elementary particle. This problem has been discussed by several authors.
@@
As noted in Sec. 1, at the quantum level, Poincare symmetry is a good approximate symmetry if the eigenvalues of the operators $M_{4\mu}$ are much greater than the eigenvalues of the operators $M_{\mu\nu}$ ($\mu,\nu=0,1,2,3$). As shown in Sec. 4, for a single supersingleton there is no scenario when these conditions are met but an explicit solution with such properties exists within the framework of the approach where standard elementary particles are bound states of two Dirac supersingletons such that their states satisfy the conditions (11). The problem that needs to be investigated is the accuracy with which these conditions must be met for existing elementary particles.
Acknowledgments. The author is grateful to Vladimir Karmanov for useful remarks.

### v2

@@ 6 Conclusion
As shown in the famous Dyson’s paper [3], even from purely mathematical considerations (without any physics) it follows that Poincare quantum symmetry is a special degenerate case of de Sitter quantum symmetries. As shown in [2, 4], in standard quantum theory (over the field of complex numbers), dS symmetry is more general than AdS one but in the framework of dS symmetry it is not possible to describe neutral elementary particles, i.e., particles which are equivalent to the their antiparticles. In particular, even the photon cannot be a neutral elementary particle. This problem has been discussed by several authors.
@@
As noted in Sec. 1, at the quantum level, Poincare symmetry is a good approximate symmetry if the eigenvalues of the operators $M_{4\mu}$ are much greater than the eigenvalues of the operators $M_{\mu\nu}$ ($\mu,\nu=0,1,2,3$). As shown in Sec. 5, for a single supersingleton there is no scenario when these conditions are met but explicit solutions with such properties exist when:
  • a particle which in the standard theory is considered a neutral elementary particle consists of a supersingleton and its antiparticle, and the result follows from Eqs. (8,11,12).
  • a particle which in the standard theory is considered elementary, consists of two supersingletons satisfying the semiclassical approximation, and their states satisfy the conditions (17).
Acknowledgments. The author is grateful to Vladimir Karmanov for useful discussions.